NITRIC ACID: ABSOLUTE INTENSITIES OF BAND AND RO-VIBRATIONAL LINE PROFILES

CHARLES CHACKERIAN, NASA Ames Research Center and SETI Institute, Mail Stop 245-5, Moffett Field, CA 94035-1000; STEVEN W. SHARPE, \underline{THOMAS A. BLAKE}, ROBERT L. SAMS, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC06-76RLO 1830).

Moderate-resolution spectra (0.112 cm^{−1}) (MRS) and high-resolution spectra (0.002 − 0.003 cm^{−1}) (HRS) of anhydrous nitric acid have been recorded between about 820 cm^{−1}and 3600 cm^{−1}. The MRS were recorded using a Bruker IFS 66V FTIR for at least five pressure burdens at each of three temperatures (278, 298, 323 K) in a temperature regulated cell with BaF_{2} windows. Nominal pathlength was 20 cm. The static fills of anhydrous nitric acid samples were back filled with nitrogen to a total pressure of 760 ± 5 Torr. Small absorbances from H_{2}O, CO_{2}, NO, NO_{2} and N_{2}O were accounted for and subtracted from the spectra. The corrected spectra were used to obtain integrated band absorption coefficients as well as a compilation of absorption coefficients tabulated at 0.056 cm^{−1}intervals. Within an experimental error of about two percent, the integrated band intensities are independent of temperature and in good agreement with results obtained by Giver et al.^{a} HRS of nitric acid were recorded with a Bruker IFS 120HR FTIR using a flow of neat anhydrous nitric acid through a temperature regulated (296 K) cell with a nominal pathlength of 10 cm and BaF_{2} windows. At least four pressure burdens were measured for each spectral region. Progress will be reported on the line by line analysis of the HRS for intensities in the 845 − 945 cm^{−1}(ν_{5}, 2ν_{9}), 1270 − 1370 cm^{−1}(ν_{3}, ν_{4}), and 1660 − 1760 cm^{−1}(ν_{2}) spectral regions.

aL. P. Giver, F. P. J. Valero, D. Goorvitch, F. S. Bonomo, J. Opt. Soc. Am. B1, 715 (1984).